Circular No. 7681

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001dn IN NGC 662 G. M. Hurst, Basingstoke, England, reports the discovery by T. Boles

(Coddenham, Suffolk) of an apparent supernova (mag 15.5) on an unfiltered CCD image (limiting mag 18.5) obtained on Aug. 14.069 UT with a 0.36m reflector. The following averaged astrometry was measured from six exposures: α = 1^{h}44^{m}35^{s}.32, δ = +37^{o}41′49′′.7 (equinox 2000.0), which is 4′′ west and 3′′ north of the center of NGC 662. SN 2001dn is also present on unfiltered CCD images taken on Aug. 14.870 (by Boles; mag ∼ 15.5) and 15.003 (by M. Armstrong, Rolvenden, Kent). SN 2001dn was not present on a Boles image taken on 1999 Oct. 16 (limiting mag 18.5) or on Palomar Sky Survey photographs taken on 1989 Nov. 18 and 1988 Sept. 13.

COMET C/2001 A2 (LINEAR) P. D. Feldman, H. A. Weaver, and E. B. Burgh, Johns Hopkins Uni-

versity, report observations of comet C/2001 A2 with the Far Ultraviolet Spectroscopic Explorer beginning July 12.58 UT and coinciding with the photometric event reported on IAUC 7679: “Spectra (range 91–118 nm; spectral resolution 0.03 nm) were obtained using the 30′′ × 30′′ aperture. Several new cometary emissions were identified, particularly the (0,0) bands of the CO Birge-Hopfield systems (C–X and B–X) at 108.8 and 115.1 nm, respectively; O I (^{1}D–^{1}D) at 115.2 nm; and three lines of the H_{2} Lyman system at 107.16, 111.86, and 116.68 nm, pumped by solar Lyβ fluorescence. Also detected were O I multiplets at 98.9, 102.7, and 104.0 nm, and several lines of the H I Lyman series. The rotational envelopes of the CO bands are resolved and appear to consist of both cold and warm components, the cold component accounting for 80 percent of the flux and having a rotational temperature of 60 K. The warm component may be indicative of a CO_{2} source. Both the CO bands and the O I 115.2-nm emission (an indicator of H_{2}O production) decreased by a factor of two over the 7.5 hr observation. Preliminary estimates of the production rates at the beginning of the observation are Q(CO) = 4 × 10^{27} s^{−1} and Q(H_{2}O) = 3 × 10^{29} s^{−1} (vectorial model). These values may be uncertain by as much as a factor of two, due to uncertainties in the solar flux. No emission is detected from Ar I at 104.8 and 106.7 nm and He I at 58.4 nm (in second order). We derive Q(Ar) ≤ 6 × 10^{25} s^{−1} (5σ upper limit), which implies that Ar/O is more than a factor of ten less than solar. In addition to the features listed above, more than two dozen other emissions remain unidentified.”

2001 August 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green